Briana wants to find the height of a tree in her backyard but she is afraid of heights, she thinks she can find the height of the tree by using similar triangles and stands 95ft from the base of the tree so the tip of her shadow lines of from the shadow of the tree briana is 66in tall and her shadow is 7ft long how does briana know that the two triangles are similar and what is the height of the tree
Answer by harpazo(655)

Before solving the problem, we must convert 66 inches to feet.


There are 12 inches in a foot.


So, 66/12 = 5.5 feet.

The needed proportion is 5.5/h = 7/102

Let h be the height of the tree.

(5.5)(102) = 7h

561 = 7h

561/7 = h

80.142857143 = h

We can round this decimal to the nearest unit

and the tree is about 80 feet tall.

Understand?